OpenStudy (anonymous):

A piece of wire 7 m long is cut into two pieces. One piece is bent into a square and the other is bent into an equilateral triangle. How much wire should be used for the square in order to minimize the total area? I have already tried lots of answers but none of them are right.. only one chance left..!!!!

OpenStudy (anonymous):

i am going to guess at 0 for the square, but let me do some calculation since you only have one chance

OpenStudy (anonymous):

thankkkksss !!!

OpenStudy (anonymous):

hmm maybe i am wrong. take say \(x\) for the triangle, then the side of the triangle is \(\frac{x}{3}\) and the area is \(\frac{\sqrt{3}}{4}(\frac{x}{3})^2\)

OpenStudy (anonymous):

that leaves \(7-x\) for the square, cut that in 4 parts and get an area of \((\frac{7-x}{4})^2\)

OpenStudy (anonymous):

@Mertsj if i am messing up let me know

OpenStudy (anonymous):

now we want to minimize \[A=\frac{\sqrt{3}}{4}(\frac{x}{3})^2+(\frac{7-x}{4})^2\]

OpenStudy (anonymous):

i would have thought we used it all for the triangle, but that is not the result i am getting. am i doing something wrong? anyone?

OpenStudy (mertsj):

I'm with you so far. Same thing I got. Just finish up. I got 3.04 for square

OpenStudy (anonymous):

i think u r right -- for the expression of A

OpenStudy (anonymous):

\[\frac{\sqrt{3}x^2}{36}+\frac{(x-7)^2}{16}\] take derivative set equal zero and solve or else since this is a poly of degree 2, expand this mess and find the vertex

OpenStudy (anonymous):

oh damn damn damn you want "how much for the square"!! so should have made \(x\) the amount used for the square

OpenStudy (mertsj):

No. Easier the way you did it. Just get answer and subtract from 7

OpenStudy (anonymous):

then \[A=\frac{x^2}{4}=\frac{\sqrt{3}(7-x)^2}{36}\]

OpenStudy (anonymous):

yeah i suppose you are right

OpenStudy (anonymous):

\(7-3.95525=3.04475\)

OpenStudy (anonymous):

here of course i cheat because i will be damned if i am going to try this will pencil and paper http://www.wolframalpha.com/input/?i=sqrt%283%29x^2%2F36%2B%287-x%29^2%2F16

OpenStudy (mertsj):

And if we pay attention to significant digits then 4 inches for triangle and 3 inches for square.

OpenStudy (anonymous):

guess my intuition was totally wrong. i would have said minimize by using all on the triangle

OpenStudy (anonymous):

hold on A=x^2/4 =[(3)^(1/2)(7-x)^2]/36 why?

OpenStudy (anonymous):

the answer is correct !!!!!!!!

OpenStudy (mertsj):

Answers by Satellite...never fail.

OpenStudy (anonymous):

thank you guys ! I follow you before you say "Then A=... "

OpenStudy (anonymous):

@Mertsj haha :D